A particle moves along the x-direction with constant acceleration. The displacemnt, measured from a convenient position, is 2m at time t=0 and is zero when t=10sec. If vel. of the particle is momentary zero when t=6sec, determine the acceleration a and velocity v when t=10sec.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

let convinient point is origin

 and initial velocity is u  in positive x direction

and acceleration is a in positive x direction

so

S= ut +1/2 * a* t2

-2 = u*10 +1/2  *a*100

-1/5 =u + 5a ....................1

 

and v=u+at

     0 =u+a*6

   u=-6a .......................2

 

solve equation 1 and 2

a=1/5

u=-6/5

so - sign tells that initial velocity is in the negative x dorection.

accleratin at time t = 1/5 m/sec2

velocity  v= u +at

               =-6/5 + 1/5 *10

              =4/5 m/sec

  

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 

 All the best.
 
Regards,
Askiitians Experts
Badiuddin